Summer School

Geometry of quiver-representations and preprojective algebras^{1}

10-17 September 2000 at Isle of Thorns (Sussex, UK)

Programme

Sunday 10 September (afternoon) *** Arrival

Monday 11 September (morning)

8.45-8.55 Opening
9.00-9.50 R. Rouquier

Affine quotients

10.05-10.55 Antwerp team

Representation spaces of quivers with relations Coffee

11.30-12.20 G. Bobinski, C. Czanto

Degenerations of modules correspond to extensions

Monday, 11 September (afternoon)

15.30-16.20 R.K. Sethi, J. E. Vatne

Stable and semi-stable points

16.35-17.25 K. Baur, S. Mohrdieck

Moduli spaces

17.40-18.30 D. Rumynin, D. Sheiham

Universal bundles

Tuesday, 12 September (morning)

9.00-9.50 A-M Aubert, E. Goddelle, A. Zimmermann

General representations of quivers

10.05-10.55 B.T. Jensen, D. Madsen

Maps between general representations Coffee

11.30-12.20 J. Bender, C. Chesne

Schubert varieties

Tuesday, 12 September (afternoon)

15.30-16.20 I. Gordon, H. Nagase

Schubert calculus

16.35-17.25 A. Cox, M. Fayers, K. Rietsch

Chern class calculations

17.40-18.30 Bucharest team

Introduction to Kac-Moody algebras

1www.amsta.leeds.ac.uk/∼pmtwc/quiver/.

Supported by European Commission (TMR-Scheme), EPSRC, LMS, and SFB 343 (Bielefeld).

Wednesday, 13 September (morning)

9.00-9.50 D. Kussin, F. Stumbo

Kac’s Theorem on indecomposable reps of quivers

10.05-10.55 R. Bocklandt, B. Sevenhant

Symplectic forms & preprojective algebra Coffee

11.30-12.20 A, Hubery, J. Schroer

Nilpotent Variety

Wednesday 13 September (afternoon) *** Discussion

Thursday, 14 September (morning)

9.00-9.50 G. Carnovale, R. Marsh, J. Shepherd

Crystal bases in quantized enveloping algebras

10.05-10.55 O. Khomenko

Geometric construction of crystal bases Coffee

11.30-12.20 L. Hille

Construction of the positive part for Dynkin Lie algebras

Thursday, 14 September (afternoon)

15.30-16.20 M. Reineke

Constructible functions and enveloping algebras

16.35-17.25 A. Maffei

Nakajima’s quiver varieties

17.40-18.30 O. Schiffmann

Construction of integrable representations

Friday, 15 September (morning)

9.00-9.50 V. Ginzburg

Non-commutative Geometry, Wilson’s Grassmannian and Quiver Varieties, I

10.05-10.55 G. Zwara

Types of singularities of orbit closures in module varieties Coffee

11.30-12.20 L. Le Bruyn

Noncommutative Geometry (and beyond), I

Friday, 15 September (afternoon)

15.30-16.20 H. Nakajima

Quiver varieties and quantum affine algebras, I

16.35-17.25 A. Schofield

The birational classification of moduli spaces, I

17.40-18.30 H. Derksen

On the σ-stable decomposition

Saturday, 16 September (morning)

9.00-9.50 G. Lusztig

Constructible functions on varieties attached to quivers, I

10.05-10.55 L. Le Bruyn

Noncommutative Geometry (and beyond), I Coffee

11.30-12.20 V. Ginzburg

Non-commutative Geometry, Wilson’s Grassmannian and Quiver Varieties, II

Saturday, 16 September (afternoon)

15.30-16.20 H. Nakajima

Quiver varieties and quantum affine algebras, II

16.35-17.25 A. Schofield

The birational classification of moduli spaces, II

17.40-18.30 G. Lusztig

Constructible functions on varieties attached to quivers, II

Sunday, 17 September (morning)

9.00-9.50 G. Lusztig

Constructible functions on varieties attached to quivers, III

10.05-10.55 H. Nakajima

Quiver varieties and quantum affine algebras, III Coffee

11.30-12.20 V. Ginzburg

Non-commutative Geometry, Wilson’s Grassmannian and Quiver Varieties, III

Sunday 17 September (afternoon) *** Departure